# Controlling the quantum phase transition in a double quantum dot Josephson junction via interactions

### v1

@@ III.2 QPT are induced by varying the interaction parameters
Figure 4(a) displays the entropy $S$ as a function of QD1’s interaction strength $U_{1}$ at a low temperature $T=0.005$, with fixed parameters $U_{2}=10\Delta$ and $U_{12}=0$ for QD2 and interdot interaction respectively. The entropy undergoes a sharp transition at $U_{1}\approx 1.6\Delta$, where it drops from $S=\ln 2$ (indicating two-fold degenerate ground states) to $S=0$ (signaling a non-degenerate singlet ground state), clearly marking the QPT. This transition is further characterized in Fig. 4(b) through the parity $P$, which changes from odd ($P=\pm 1$) in the pre-QPT phase to even ($P=0$) in the post-QPT phase. The spin-spin correlation between dots, shown in Fig. 4(c), reveals the formation of an antiferromagnetic state after the QPT, while remaining nearly uncorrelated before the transition. Finally, Fig. 4(d) demonstrates the superconducting order parameters $\langle d_{1\downarrow}d_{1\uparrow}\rangle$ and $\langle d_{2\downarrow}d_{2\uparrow}\rangle$ as functions of $U_{1}$: in the pre-QPT phase ($U_{1}<1.6\Delta$), QD1 maintains significant pairing ($\langle d_{1\downarrow}d_{1\uparrow}\rangle\gg 0$) characteristic of an ABS while QD2 shows negligible pairing ($\langle d_{2\downarrow}d_{2\uparrow}\rangle\approx 0$) corresponding to a single-electron state; post-QPT ($U_{1}>1.6\Delta$), both order parameters vanish ($\langle d_{1\downarrow}d_{1\uparrow}\rangle,\langle d_{2\downarrow}d_{2\uparrow}\rangle\rightarrow 0$), indicating the complete suppression of superconductivity and transition to single-electron states in both quantum dots. Remarkably, the two quantum dots establish a non-local Cooper pairing under these specific conditions.
@@
For Fig. 4, our analysis of the pre-transition ground state has been comprehensively presented, while the post-transition ground state can be unambiguously determined from the previously examined physical quantities to be the antiferromagnetic singlet state $|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle$, which exhibits zero parity ($P=0$) and features both QD1 and QD2 in single-particle states yet quantum-mechanically coupled through the formation of an interdot Cooper pair, as evidenced by the finite spin-spin correlation $\langle\vec{S}{1}\!\!\cdot\vec{S}{2}\rangle<0$ and the complete suppression of local superconducting order parameters $\langle d_{i\downarrow}d_{i\uparrow}\rangle=0$ ($i=1,2$) in the post-transition regime. At the particle-hole symmetric point, the Coulomb interaction effectively suppresses both doubly occupied and empty states, thereby enforcing QD1 and QD2 into single-electron configurations; however, the superconducting proximity effect mediates the formation of a nonlocal Cooper pair between the quantum dots, resulting in the post-QPT ground state being characterized by the antiferromagnetic singlet state $(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)/\sqrt{2}$ that combines single-electron localization with nonlocal superconducting correlations.
@@
In Fig. 4(f), we systematically examine the Josephson current $I_{J}$ as a function of the superconducting phase difference $\phi$ for varying interaction strengths $U_{1}$, where we observe distinct current-phase relations that correspond to different quantum states: for the non-interacting case ($U_{1}=0$), the current exhibits characteristic $0^{\prime}(\pi^{\prime})$-phase behavior, while at intermediate interaction strength ($U_{1}=3\Delta$), the current-phase relation undergoes a $\pi$-shift that reflects the system’s transition to a $\pi$-phase state, a direct consequence of QD1 entering a single-electron configuration where the superconducting correlations vanish ($\langle d_{1\downarrow}d_{1\uparrow}\rangle=0$) and single-particle tunneling processes dominate the transport.
Fig. 5: (Colour online) (a) The interdot spin correlation $\langle\vec{S}{1}\!\!\cdot\!\!\vec{S}{2}\rangle$ as a function of the uniform Coulomb interaction strength $U=U_{1}=U_{2}$ in both quantum dots. (b) Phase diagram characterizing the singlet-doublet state transition as a function of the normalized interaction parameters $U_{1}/\Delta$ and $U_{2}/\Delta$, with all other system parameters maintained identical to those specified in Fig. 3.
Our earlier calculations first introduced electron-electron interaction specifically on QD2, revealing a QPT, followed by introducing interaction on QD1 while maintaining finite interaction on QD2, which induced another QPT; nevertheless, the fundamental question of whether similar phase transitions would emerge when interactions are introduced first on QD1 or simultaneously on both QDs remained unresolved, prompting the systematic investigation presented in Fig. 5(a) where we set $U_{1}=U_{2}=U$ and examined the spin correlation function $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ as a continuous function of $U$, finding smooth evolution without any non-analytic behavior that would signal a QPT - in this symmetric interaction case, the ground state evolves smoothly from the non-interacting configuration $A_{1}|0_{1}0_{2}\rangle+B_{1}|0_{1}d_{2}\rangle+C_{1}|d_{1}0_{2}\rangle+D_{1}|d_{1}d_{2}\rangle+E_{1}(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)$ to the strongly correlated limit $(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)/\sqrt{2}$, as the increasing Coulomb interactions $U$ suppress both doubly occupied and empty states at the particle-hole symmetric point, causing the coefficients $A_{1}$, $B_{1}$, $C_{1}$ and $D_{1}$ to diminish while $E_{1}$ remains dominant, with this comprehensive picture being further corroborated by the phase diagram in Fig. 5(b) that maps the singlet-doublet phase boundary across the full parameter space of $U_{1}$ and $U_{2}$ interactions.
@@
Figure 6(a) displays the entropy $S$ versus the interdot Coulomb interaction $U_{12}$ at temperature $T=0.005$, with fixed on-site interactions $U_{1}=U_{2}=10\Delta$, revealing a QPT at $U_{12}\approx 4.3\Delta$ where the ground state changes from a singlet (pre-QPT) to a doublet (post-QPT), as further evidenced by the parity $P$ in Fig. 6(b) transitioning from 00 to ±$\pm 1$ across the QPT point, while Fig. 6(c) shows the interdot spin correlation $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ maintaining antiferromagnetic character ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle<0$) before the QPT but vanishing ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle\approx 0$) afterward, and Fig. 6(d) demonstrates the occupation numbers $n_{1}$ and $n_{2}$ of both quantum dots deviating from unity at finite $U_{12}$ even at the nominal particle-hole symmetric point $\epsilon_{i}=-U_{i}/2$ due to $U_{12}$-induced particle-hole symmetry breaking, with both occupations exhibiting an abrupt decrease at the QPT point that coincides with the entropy, parity, and spin correlation transitions.
@@ III.3 Ferromagnetic-antiferromagnetic phase transition
Our previous analysis has demonstrated that the system enters an antiferromagnetic state with negative spin correlation when both QD1 and QD2 experience strong interdot Coulomb interaction, though these results were obtained in the zero-field limit where the Zeeman effect is absent; the introduction of an external magnetic field $B$ may potentially induce a transition to a ferromagnetic configuration ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle>0$) through spin polarization, which motivates our systematic investigation in Fig. 7 of the magnetic field dependence of the quantum phase transitions.
Fig. 7: (Colour online) (a) $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ as a function of $U_{1}$ for various external magnetic field. (b) The spin correlation as a function of external parallel magnetic field $B=-B_{1}=-B_{2}$ with $U_{1}=10\Delta$. (c) $I_{J}/I_{0}$ as a function of external parallel magnetic field when phase difference between two SC leads is $0.3\pi$. The interaction of QD2 is $U_{2}=10\Delta$. The other parameters are given the same as in Fig. 3 unless specifically stated.
Figure 7(a) systematically examines the interdot spin correlation as a function of $U_{1}$ under various magnetic field configurations while maintaining $U_{2}=10\Delta$ and $U{12}=0$, demonstrating that a magnetic field applied solely to QD1 reduces the critical interaction strength $U_{1c}$ for the QPT by disrupting the Andreev bound state (ABS) in QD1 through Zeeman splitting that effectively mimics additional Coulomb interaction, whereas applying the field exclusively to QD2 increases $U_{1c}$ since the magnetic field’s influence on QD2 is energetically equivalent to enhancing its local interaction strength, as corroborated by the phase diagram in Fig. 5(b) showing expanded parameter regions for the singlet phase; when co-directional magnetic fields $B_{1}=B_{2}=B$ are applied to both QDs, the post-QPT ground state transitions from antiferromagnetic ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle<0$) to ferromagnetic ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle>0$) ordering, though the precise nature of this transition under strong interactions ($U_{1}=U_{2}=10\Delta$) remained ambiguous until Fig. 7(b) revealed an abrupt change in spin correlations at a critical field strength $B_{c}\approx 0.2$, unambiguously demonstrating the magnetic field’s capability to induce a sharp QPT between these competing magnetic phases when $U_{12}=0$.
@@ III.4 Nonlocal magnetization
Our prior investigations have rigorously demonstrated that the system exhibits odd parity ($P=\pm 1$) under two distinct interaction regimes: (i) the $U_{1}=U_{12}=0$ with strong $U_{2}$ coupling, and (ii) the symmetric strong-coupling scenario where $U_{1}=U_{2}=10\Delta$ with large interdot interaction $U_{12}$, with the application of a weak magnetic field $B_{1}=-0.001$ to QD1 inducing Zeeman splitting of the ground state energy levels that lifts the degeneracy and drives the system into a spin-polarized configuration with definite parity ($P=1$ or $P=-1$) and quantized total spin ($S_{z}=0.5$ or $S_{z}=-0.5$), as systematically mapped out in Fig. 8 through the detailed dependence of the individual quantum dot spins $\langle S_{1z}\rangle$ and $\langle S_{2z}\rangle$ on the interaction parameters $U_{1}$, $U_{2}$, and $U_{12}$ across both parameter regimes.
@@
Remarkably, when applying a magnetic field exclusively to QD1 (in the ABS state), the induced magnetization emerges in QD2, demonstrating that QD2’s magnetization fundamentally requires first-order electronic processes mediated through QD1. Fig. 9 delineates the two essential first-order transition mechanisms: (i) electron exchange between the superconducting reservoir and QD1, demanding $\Delta$-dependent energy to excite individual electrons, and (ii) inter-QD transitions between QD1-QD2 requiring $U_{2}/2$-scaled energy at the particle-hole symmetric point ($\varepsilon_{2\sigma}=-U_{2}/2$). The system displays two characteristic regimes: for small $U_{2}$, electron transport occurs predominantly between QD1-QD2, whereas large $U_{2}$ promotes reservoir-QD1 transitions. The spin-polarizing magnetic field on QD1 enforces spin-down polarization in QD1 following first-order transitions. In systems with small $U_{2}$ values, electron tunneling predominantly occurs between QD1 and QD2, where before tunneling QD1 acquires spin-down polarization and QD2 starts in a single-electron configuration with its spin initially pointing downward, leading to magnetization that aligns parallel with the external magnetic field. On the other hand, when $U_{2}$ is large and electrons primarily transfer between the superconducting reservoir and QD1 (while QD1 still achieves spin-down polarization and QD2 remains in its single-electron state), after tunneling the nonlocal nature of the superconducting proximity effect facilitates Cooper pair formation across QD1 and QD2, thereby reversing QD2’s spin to an upward orientation and generating magnetization that is oriented antiparallel to the applied magnetic field. The $\Delta=10$ scenario (red dashed line) validates this picture through pure QD1-QD2 transitions and resulting spin-down magnetization. Additionally, the ground state evolution from $|0_{1}\!\!\downarrow_{2}\rangle-|d_{1}\!\!\downarrow_{2}\rangle$ to $|0_{1}\!\!\uparrow_{2}\rangle-|d_{1}\!\!\uparrow_{2}\rangle$ with increasing $U_{2}$ potentially indicates a quantum phase transition.

### v2

@@ III.2 QPT are induced by varying the electron-electron interaction parameters
Figure 4(a) displays the entropy $S$ as a function of QD1’s interaction strength $U_{1}$ at a low temperature $T=0.005$, with fixed parameters $U_{2}=10\Delta$ and $U_{12}=0$ for QD2 and interdot interaction respectively. The entropy undergoes a sharp transition at $U_{1}\approx 1.6\Delta$, where it drops from $S=\ln 2$ (indicating two-fold degenerate ground states) to $S=0$ (signaling a non-degenerate singlet ground state), clearly marking the QPT. This transition is further characterized in Fig. 4(b) through the parity $P$, which changes from odd ($P=\pm 1$) in the pre-QPT phase to even ($P=0$) in the post-QPT phase. The spin-spin correlation between dots, shown in Fig. 4(c), reveals the formation of an antiferromagnetic state after the QPT, while remaining nearly uncorrelated before the transition. Finally, Fig. 4(d) demonstrates the superconducting order parameters $\langle d_{1\downarrow}d_{1\uparrow}\rangle$ and $\langle d_{2\downarrow}d_{2\uparrow}\rangle$ as functions of $U_{1}$: in the pre-QPT phase ($U_{1}<1.6\Delta$), QD1 maintains significant pairing ($|\langle d_{1\downarrow}d_{1\uparrow}\rangle|\gg 0$) characteristic of an ABS while QD2 shows negligible pairing ($\langle d_{2\downarrow}d_{2\uparrow}\rangle\approx 0$) corresponding to a single-electron state; post-QPT ($U_{1}>1.6\Delta$), both order parameters vanish ($\langle d_{1\downarrow}d_{1\uparrow}\rangle,\langle d_{2\downarrow}d_{2\uparrow}\rangle\rightarrow 0$), indicating the complete suppression of superconductivity and transition to single-electron states in both quantum dots. Remarkably, the two quantum dots establish a non-local Cooper pairing under these specific conditions.
@@
For Fig. 4, our analysis of the pre-transition ground state has been comprehensively presented, while the post-transition ground state can be unambiguously determined from the previously examined physical quantities to be the antiferromagnetic singlet state $|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle$, which exhibits zero parity ($P=0$) and features both QD1 and QD2 in single-particle states yet quantum-mechanically coupled through the formation of an interdot Cooper pair, as evidenced by the finite spin-spin correlation $\langle\vec{S}_{1}\!\!\cdot\vec{S}_{2}\rangle<0$ and the complete suppression of local superconducting order parameters $\langle d_{i\downarrow}d_{i\uparrow}\rangle=0$ ($i=1,2$) in the post-transition regime. At the particle-hole symmetric point, the Coulomb interaction effectively suppresses both doubly occupied and empty states, thereby enforcing QD1 and QD2 into single-electron configurations; however, the superconducting proximity effect mediates the formation of a nonlocal Cooper pair between the quantum dots, resulting in the post-QPT ground state being characterized by the antiferromagnetic singlet state $(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)/\sqrt{2}$ that combines single-electron localization with nonlocal superconducting correlations.
@@
In Fig. 4(f), we examine the Josephson current $I_{J}$ as a function of the superconducting phase difference $\phi$ for varying interaction strengths $U_{1}$, where we observe distinct current-phase relations that correspond to different quantum states: for the non-interacting case ($U_{1}=0$), the current exhibits characteristic $0^{\prime}(\pi^{\prime})$-phase behavior, while at intermediate interaction strength ($U_{1}=3\Delta$), the current-phase relation undergoes a $\pi$-shift that reflects the system’s transition to a $\pi$-phase state, a direct consequence of QD1 entering a single-electron configuration where the superconducting correlations vanish ($\langle d_{1\downarrow}d_{1\uparrow}\rangle=0$) and single-particle tunneling processes dominate the transport.
Fig. 5: (Colour online) (a) The interdot spin correlation $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ as a function of the uniform Coulomb interaction strength $U=U_{1}=U_{2}$ in both quantum dots. (b) Phase diagram characterizing the singlet-doublet state transition as a function of the normalized interaction parameters $U_{1}/\Delta$ and $U_{2}/\Delta$, with all other system parameters maintained identical to those specified in Fig. 3.
Our previous calculations first introduced electron-electron interaction specifically on QD2, revealing a QPT, followed by introducing interaction on QD1 while maintaining finite interaction on QD2, which induced another QPT; nevertheless, the fundamental question of whether similar phase transitions would emerge when interactions are introduced first on QD1 or simultaneously on both QDs remained unresolved, prompting the systematic investigation presented in Fig. 5(a) where we set $U_{1}=U_{2}=U$ and examined the spin correlation function $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ as a continuous function of $U$, finding smooth evolution without any non-analytic behavior that would signal a QPT - in this symmetric interaction case, the ground state evolves smoothly from the non-interacting configuration $A_{1}|0_{1}0_{2}\rangle+B_{1}|0_{1}d_{2}\rangle+C_{1}|d_{1}0_{2}\rangle+D_{1}|d_{1}d_{2}\rangle+E_{1}(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)$ to the strongly correlated limit $(|\!\!\uparrow_{1}\downarrow_{2}\rangle-|\!\!\downarrow_{1}\uparrow_{2}\rangle)/\sqrt{2}$, as the increasing Coulomb interaction $U$ suppresses both doubly occupied and empty states at the particle-hole symmetric point, causing the coefficients $A_{1}$, $B_{1}$, $C_{1}$ and $D_{1}$ to diminish while $E_{1}$ remains dominant, with this comprehensive picture being further corroborated by the phase diagram in Fig. 5(b) that maps the singlet-doublet phase boundary across the full parameter space of $U_{1}$ and $U_{2}$ interactions.
@@
Figure 6(a) displays the entropy $S$ versus the interdot Coulomb interaction $U_{12}$ at temperature $T=0.005$, with the fixed on-site interactions $U_{1}=U_{2}=10\Delta$, revealing a QPT at $U_{12}\approx 4.3\Delta$ where the ground state changes from a singlet (pre-QPT) to a doublet (post-QPT), as further evidenced by the parity $P$ in Fig. 6(b) transitioning from 00 to ±$\pm 1$ across the QPT point, while Fig. 6(c) shows the interdot spin correlation $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ maintaining antiferromagnetic character ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle<0$) before the QPT but vanishing ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle\approx 0$) afterward, and Fig. 6(d) demonstrates the occupation numbers $n_{1}$ and $n_{2}$ of both quantum dots deviating from unity at finite $U_{12}$ even at the nominal particle-hole symmetric point $\epsilon_{i}=-U_{i}/2$ due to $U_{12}$-induced particle-hole symmetry breaking, with both occupations exhibiting an abrupt decrease at the QPT point that coincides with the entropy, parity, and spin correlation transitions.
@@ III.3 Ferromagnetic-antiferromagnetic phase transition
Our previous analysis has demonstrated that the system enters an antiferromagnetic state with negative spin correlation when both QD1 and QD2 experience strong on-site Coulomb interaction, though these results were obtained in the zero-field limit where the Zeeman effect is absent; the introduction of an external magnetic field $B$ may potentially induce a transition to a ferromagnetic configuration ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle>0$) through spin polarization, which motivates our investigation in Fig. 7 of the magnetic field dependence of the quantum phase transitions.
Fig. 7: (Colour online) (a) $\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle$ as a function of $U_{1}$ for various external magnetic fields. (b) The spin correlation as a function of external parallel magnetic field $B=-B_{1}=-B_{2}$ with $U_{1}=10\Delta$. (c) $I_{J}/I_{0}$ as a function of external parallel magnetic field when phase difference between two SC leads is $0.3\pi$. The interaction of QD2 is $U_{2}=10\Delta$. The other parameters are given the same as in Fig. 3 unless specifically stated.
Figure 7(a) systematically examines the interdot spin correlation as a function of $U_{1}$ under various magnetic field configurations while maintaining $U_{2}=10\Delta$ and $U_{12}=0$, demonstrating that a magnetic field applied solely to QD1 reduces the critical interaction strength $U_{1c}$ for the QPT by disrupting the ABS in QD1 through Zeeman splitting that effectively mimics additional Coulomb interaction, whereas applying the field exclusively to QD2 increases $U_{1c}$ since the magnetic field’s influence on QD2 is energetically equivalent to enhancing its local interaction strength, as corroborated by the phase diagram in Fig. 5(b) showing expanded parameter regions for the singlet phase; when co-directional magnetic fields $B_{1}=B_{2}=B$ are applied to both QDs, the post-QPT ground state transitions from antiferromagnetic ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle<0$) to ferromagnetic ($\langle\vec{S}_{1}\!\!\cdot\!\!\vec{S}_{2}\rangle>0$) ordering, though the precise nature of this transition under strong interactions ($U_{1}=U_{2}=10\Delta$) remained ambiguous until Fig. 7(b) revealed an abrupt change in spin correlations at a critical field strength $B_{c}\approx 0.2$, unambiguously demonstrating the magnetic field’s capability to induce a sharp QPT between these competing magnetic phases when $U_{12}=0$.
@@ III.4 Nonlocal magnetization
Our prior investigations have rigorously demonstrated that the system exhibits odd parity ($P=\pm 1$) under two distinct interaction regimes: (i) the $U_{1}=U_{12}=0$ with strong $U_{2}$ coupling, and (ii) the symmetric strong-coupling scenario where $U_{1}=U_{2}=10\Delta$ with large interdot interaction $U_{12}$. The application of a weak magnetic field $B_{1}=-0.001$ to QD1 induces Zeeman splitting of the ground state energy levels, which removes the degeneracy and drives the system into a spin-polarized configuration characterized by well-defined parity ($P=1$ or $P=-1$) and quantized total spin ($S_{z}=0.5$ or $S_{z}=-0.5$).
@@
Remarkably, when applying a magnetic field exclusively to QD1 (in the ABS state), the induced magnetization emerges in QD2, demonstrating that QD2’s magnetization fundamentally requires first-order electronic processes mediated through QD1. Figure. 9 delineates the two essential first-order transition mechanisms: (a) inter-QD transitions between QD1-QD2 requiring $U_{2}/2$-scaled energy at the particle-hole symmetric point ($\varepsilon_{2\sigma}=-U_{2}/2$), and (b) electron exchange between the superconducting reservoir and QD1, demanding $\Delta$-dependent energy to excite individual electrons. The system displays two characteristic regimes: for small $U_{2}$, electron transport occurs predominantly between QD1-QD2, whereas large $U_{2}$ promotes reservoir-QD1 transitions. The spin-polarizing magnetic field on QD1 enforces spin-down polarization in QD1 following first-order transitions. In systems with small $U_{2}$ values, electron tunneling predominantly occurs between QD1 and QD2, where before tunneling QD1 acquires spin-down polarization and QD2 starts in a single-electron configuration with its spin initially pointing downward, leading to magnetization that aligns parallel with the external magnetic field. On the other hand, when $U_{2}$ is large and electrons primarily transfer between the superconducting reservoir and QD1 (while QD1 still achieves spin-down polarization and QD2 remains in its single-electron state), after tunneling the nonlocal nature of the superconducting proximity effect facilitates Cooper pair formation across QD1 and QD2, thereby reversing QD2’s spin to an upward orientation and generating magnetization that is oriented antiparallel to the applied magnetic field. The $\Delta=10$ scenario (red dashed line) validates this picture through pure QD1-QD2 transitions and resulting spin-down magnetization. Additionally, the ground state evolution from $|0_{1}\!\!\downarrow_{2}\rangle-|d_{1}\!\!\downarrow_{2}\rangle$ to $|0_{1}\!\!\uparrow_{2}\rangle-|d_{1}\!\!\uparrow_{2}\rangle$ with increasing $U_{2}$ potentially indicates a quantum phase transition.